# On the Domination of surface-group representations in $\mathrm{PU}(2,1)$

### v1

@@ Abstract.
This article explores surface-group representations into the complex hyperbolic group $\mathrm{PU}(2,1)$ and presents domination results for a special class of representations called $T$-bent representations. Let $S_{g,k}$ be a punctured surface of negative Euler characteristic. We prove that for a $T$-bent representation $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PU}(2,1)$, there exists a discrete and faithful representation $\rho_{0}:\pi_{1}(S_{g,k})\rightarrow\mathrm{PO}(2,1)$ that dominates $\rho$ in the Bergman translation length spectrum, while preserving the lengths of the peripheral loops.
@@ 1. Introduction
The study of surface-group representations $\rho:\pi_{1}(S_{g,k})\rightarrow G$ into different Lie groups $G$ (preferably of higher rank) is a significant aspect in Higher Teichmüller theory (see [Wie18] for a survey). In this article, we focus on representations into $\mathrm{PU}(2,1)$. Recall that, $\mathrm{PU}(2,1)$ is the group of holomorphic isometries of 2 dimensional complex hyperbolic plane ${\mathbf{H}}^{2}_{\mathbb{C}}$. Despite sharing some structural similarities with their real counterpart, surface-group representations into $\mathrm{PU}(2,1)$ exhibit distinct geometric and dynamical behaviors that are still not fully understood. Our aim is to explore these representations in relation to domination, a notion that has been used in analyzing surface-group representations in certain other settings.
@@
Domination in the context of surface-group representations first arose in the context of anti-de Sitter (AdS) $3$-manifolds, cf. [KP94]. In 2015, Guéritaud-Kassel-Wolff [GKW15] showed that for a closed surface $S_{g}$ and a non-Fuchsian representation $\rho:\pi_{1}(S_{g})\rightarrow\mathrm{PSL}_{2}(\mathbb{R})$, there exists a Fuchsian representation $j:\pi_{1}(S_{g})\rightarrow\mathrm{PSL}_{2}(\mathbb{R})$ that strictly dominates $\rho$. This theorem has direct application in the construction of compact AdS 3-manifolds. Around the same period, Deroin-Tholozan [DT15] showed that a representation $\rho:\pi_{1}(S_{g})\rightarrow\mathrm{PSL}_{2}(\mathbb{C})$ can be strictly dominated by a Fuchsian representation $\rho_{0}:\pi_{1}(S_{g})\rightarrow\mathrm{PSL}_{2}(\mathbb{R})$, unless $\rho$ is itself Fuchsian. Other progresses in this area can also be found in [DMSV19], [Sag23], [Tho17].
For a punctured surface $S_{g,k}$, Gupta-Su [GS23] showed that for a non-Fuchsian representation $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PSL}_{2}(\mathbb{C})$ with $n\geq 1$, there exists a Fuchsian representation $\rho_{0}:\pi_{1}(S_{g,k})\rightarrow\mathrm{PSL}_{2}(\mathbb{R})$ that strictly dominates $\rho$, preserving the lengths of the peripheral loops. Note that, if we allow the lengths of the peripheral curves to increase also, then such a domination can be easily constructed by the strip deformation, as discussed in [GS23]. In [BG25], their result has been generalized for the Hilbert length spectrum and the translation length spectrum to the case when the target-group is $\mathrm{PSL}_{n}(\mathbb{C})$.
While this notion of domination has been studied in some depth for target groups like $\mathrm{PSL}_{n}(\mathbb{C})$, $n\geq 2$, much less is known when the target is $\mathrm{PU}(2,1)$. We shall explore this in this paper. For this case, we turn to a specific kind of representations, namely, $T$-bent representations, introduced by Will [Wil12]. These representations are constructed using a fixed ideal triangulation of the surface and a certain boundary map.
@@
The surface-group representation variety $\text{Rep}_{\pi_{1}(S_{g,k}),\mathrm{PU}(2,1)}$ has real dimension $16g-16+8n$. However, Will showed that $\mathcal{BR}_{T}$ has real dimension $12g-12+6k$ (see 2.7).
In this article, we show domination for these $T$-bent representations into $\mathrm{PU}(2,1)$. Let $\sigma(A)$ denote the spectral radius of $A$. Recall that,
@@
Let $\ell(A)$ denote the Bergman translation length (see equation 2.1 for the metric) of an element $A\in\mathrm{SU}(2,1)$. From Proposition 3.10 in [Par12], we know that $\ell(A)=2\ln\sigma(A).$ For an element $A\in\mathrm{PU}(2,1)$, we take a representative $\tilde{A}\in\mathrm{U}(2,1)$ and define
$\ell(A):=2\ln\sigma(\tilde{A}).$
Note that any two lifts of $A$ to $\mathrm{U}(2,1)$ differ by multiplication with a scalar $\lambda$ such that $\lambda^{3}=1$. So they have the same spectral radius. Consequently, Bergman translation length is well-defined on $\mathrm{PU}(2,1)$.
@@
for all $\gamma\in\pi_{1}(S_{g,k})$.
Here we prove the following.
@@
Let $S_{g,k}$ be an oriented surface of negative Euler characteristic with at least one puncture. For a $T$-bent representation $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PU}(2,1)$, there exists a discrete and faithful representation $\rho_{0}:\pi_{1}(S_{g,k})\rightarrow\mathrm{PO}(2,1)$ that dominates $\rho$. Moreover, if $\gamma\in\pi_{1}(S_{g,k})$ is a peripheral loop, then the Bergman translation length of $\rho(\gamma)$ remain unchanged.
We have used the ${\tt Z}$-invariant (see Section 2.3) associated with a pair of adjacent real ideal triangles to establish our result. Using these coordinates, we can similarly define the bending fiber of a $T$-bent representation $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PU}(2,1)$ as in [BG25, Definition 2.14]. So 1.3 can also be restated as:
@@
The trace of an element $A\in\mathrm{SU}(2,1)$ also determines its geometric action and is directly associated with $\ell(A)$, as described in [Par12]. As an immediate corollary of 1.3, we also get:
@@
$|tr(\rho(\gamma))|\leq|tr(\rho_{0}(\gamma))|$
@@
We have also experimentally observed that the inequality does not extend for the discriminator function $f(\mathrm{tr}(\rho(\gamma)))$ (see 2.1 and the Appendix).
@@ Acknowledgements
The authors would like to thank Subhojoy Gupta for indicating this problem to us and for motivations and useful discussions. The authors also thank John R. Parker for valuable discussions.
@@ 2.1. Complex hyperbolic plane and its isometries
[…]
$\langle\textbf{z},\textbf{w}\rangle_{1}=z_{1}\bar{w}_{1}+z_{2}\bar{w}_{2}-z_{3}\bar{w}_{3}\ \text{ and }\ \langle\textbf{z},\textbf{w}\rangle_{2}=z_{1}\bar{w}_{3}+z_{2}\bar{w}_{2}+z_{3}\bar{w}_{1}.$
These are given by the Hermitian matrices
$J_{1}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix}\ \text{ and }\ J_{2}=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$
respectively, i.e., $\langle\textbf{z},\textbf{w}\rangle_{m}=\textbf{w}^{*}J_{m}\textbf{z}$ for $m=1,2$. Unless otherwise specified, we will work with the second Hermitian form in this article.
The Heisenberg coordinates (see [Par04, §4] or [Wil12, §2] for more details) of a point $(z_{1},z_{2},1)^{t}\in\partial{\mathbf{H}}^{2}_{\mathbb{C}}$ is given by $[\zeta,t]$, where $z_{2}=\zeta\sqrt{2}$ and $z_{1}=-|\zeta|^{2}+it$.
@@
The holomorphic isometry group of ${\mathbf{H}}^{2}_{\mathbb{C}}$ is the projective unitary group $\mathrm{PU}(2,1)$, consisting of projective linear transformations that preserve the Hermitian form (up to scalar multiple). Explicitly, $\mathrm{PU}(2,1)=\mathrm{U}(2,1)/\{\lambda I\}$, where $\mathrm{U}(2,1)$ is the group of matrices $A\in\mathrm{GL}_{3}(\mathbb{C})$ satisfying $A^{*}JA=J$, with $J$ being the Hermitian matrix corresponding to the chosen form and $\lambda\in\mathbb{C}^{*}$. This group acts transitively and holomorphically on ${\mathbf{H}}^{2}_{\mathbb{C}}$, preserving both the complex structure and the Bergman metric. The isometry group Isom(${\mathbf{H}}^{2}_{\mathbb{C}}$) is generated by $\mathrm{PU}(2,1)$ and the complex conjugation (antiholomorphic). Throughout this article, we will primarily consider holomorphic isometries (elements of $\mathrm{PU}(2,1)$), unless stated otherwise.
@@
[Par04, Theorem 3.17] Let $f(z)=|z|^{4}-8\Re(z^{3})+18|z|^{2}-27$. For $A\in\mathrm{SU}(2,1)$, the following holds:
  • $A$ has an eigenvalue $\lambda$ with $|\lambda|\neq 1$ if and only if $f(\mathrm{tr}(A))>0$; in this case, $A$ is loxodromic,
  • $A$ has a repeated eigenvalue if and only if $f(\mathrm{tr}(A))=0$; in this case, $A$ is parabolic,
  • $A$ has distinct eigenvalues all of unit modulus if and only if $f(\mathrm{tr}(A))<0$; in this case, $A$ is elliptic.
@@ 2.2. Totally geodesic subspaces and ideal triangles
  • Complex lines, which are biholomorphic to the Poincaré disk, and thus an embedded copy of ${\mathbf{H}}^{1}_{\mathbb{C}}$ inside ${\mathbf{H}}^{2}_{\mathbb{C}}$. Each complex line is the intersection of ${\mathbf{H}}^{2}_{\mathbb{C}}$ with a complex projective line lying entirely in the negative cone.
  • Totally real Lagrangian planes, (also called real planes) which are isometric copies of the Klein-Beltrami model of real hyperbolic plane $\mathbb{H}^{2}_{\mathbb{R}}$. These arise as the intersection of ${\mathbf{H}}^{2}_{\mathbb{C}}$ with totally real 2-planes in $\mathbb{C}^{3}$.
@@
[Gol99, §7] Let $x=(x_{1},x_{2},x_{3})$ and $y=(y_{1},y_{2},y_{3})$ be two ideal triangles. Then there is a holomorphic (resp. antiholomorphic) isometry $A\in$ Isom $({\mathbf{H}}^{2}_{\mathbb{C}})$ with $A(x_{i})=y_{i}$ if and only if they have same (resp. opposite) Cartan invariant. Moreover, an ideal triangle has Cartan invariant 0 (resp. ±$\pm\pi/2$) if and only if it is contained in a real plain (resp. complex line).

### v2

@@ Abstract.
This article explores surface-group representations into the complex hyperbolic isometry group $\mathrm{PU}(2,1)$ and presents domination results for a special class of representations called the $T$-bent representations. Let $S_{g,k}$ be a surface of genus $g$ with $k>0$ punctures and negative Euler characteristic. We prove that for a $T$-bent representation $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PU}(2,1)$, there exists a discrete and faithful representation $\rho_{0}:\pi_{1}(S_{g,k})\rightarrow\mathrm{PO}(2,1)$ that dominates $\rho$ in the Bergman translation length spectrum, while preserving the lengths of the peripheral loops.
@@ 1. Introduction
The study of surface-group representations $\rho:\pi_{1}(S_{g,k})\rightarrow G$ into different Lie groups $G$ (preferably of higher rank) is a significant aspect in Higher Teichmüller theory (see [Wie18] for a survey). In this article, we focus on representations into $\mathrm{PU}(2,1)$. Recall that, $\mathrm{PU}(2,1)$ is the group of holomorphic isometries of the two dimensional complex hyperbolic plane ${\mathbf{H}}^{2}_{\mathbb{C}}$. Despite sharing some structural similarities with their real counterpart, surface-group representations into $\mathrm{PU}(2,1)$ exhibit distinct geometric and dynamical behaviors that are still not fully understood. Our aim is to explore these representations in relation to domination, a notion that has been used in analyzing surface-group representations in certain other settings.
@@
Domination in the context of surface-group representations first arose in the study of anti-de Sitter (AdS) $3$-manifolds, cf. [KP94]. It is an important geometric phenomenon with significant implications for surface-group representations. In 2015, Guéritaud-Kassel-Wolff [GKW15] showed that for a closed surface $S_{g}$ and a non-Fuchsian representation $\rho:\pi_{1}(S_{g})\rightarrow\mathrm{PSL}_{2}(\mathbb{R})$, there exists a Fuchsian representation $j:\pi_{1}(S_{g})\rightarrow\mathrm{PSL}_{2}(\mathbb{R})$ that strictly dominates $\rho$. This theorem has direct application in the construction of compact AdS 3-manifolds. Around the same period, Deroin-Tholozan [DT15] showed that a representation $\rho:\pi_{1}(S_{g})\rightarrow\mathrm{PSL}_{2}(\mathbb{C})$ can be strictly dominated by a Fuchsian representation $\rho_{0}:\pi_{1}(S_{g})\rightarrow\mathrm{PSL}_{2}(\mathbb{R})$, unless $\rho$ is itself Fuchsian. Other progresses in dominating surface-group representations can also be found in [Tho17], [GGKW17], [DMSV19], [Sag23].
For a punctured surface $S_{g,k}$, Gupta-Su [GS23] showed that for a non-Fuchsian representation $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PSL}_{2}(\mathbb{C})$ with $n\geq 1$, there exists a Fuchsian representation $\rho_{0}:\pi_{1}(S_{g,k})\rightarrow\mathrm{PSL}_{2}(\mathbb{R})$ that strictly dominates $\rho$, preserving the lengths of the peripheral loops. Note that, if we allow the lengths of the peripheral curves to increase also, then such a domination can be easily constructed by the ‘strip deformation’, as discussed in [GS23]. The above result has been generalized in [BG25] for the Hilbert length spectrum and the translation length spectrum to the case when the target-group is $\mathrm{PSL}_{n}(\mathbb{C})$.
While this notion of domination has been studied in some depth for target groups like $\mathrm{PSL}_{n}(\mathbb{C})$, $n\geq 2$, much less is known when the target is $\mathrm{PU}(2,1)$. We shall explore this in this paper. For this case, we turn to a specific kind of representations, namely, the $T$-bent representations, introduced by Will [Wil12]. These representations are constructed using a fixed ideal triangulation of the surface and a certain boundary map.
@@
The surface-group representation variety $\text{Rep}_{(\pi_{1}(S_{g,k}),\mathrm{PU}(2,1))}$ has real dimension 16g16+8k16g-16+8k. However, Will showed that $\mathcal{BR}_{T}$ has real dimension 12g12+6k12g-12+6k (see 2.7).
In this article, we show domination for these $T$-bent representations into $\mathrm{PU}(2,1)$. Let $\sigma(A)$ denote the spectral radius of $A$. Recall that
@@
Let $\ell(A)$ denote the translation length of an element $A\in\mathrm{SU}(2,1)$ with respect to the Bergman metric, a Riemannian metric on ${\mathbf{H}}^{2}_{\mathbb{C}}$ (see equation 2.2). From Proposition 3.10 in [Par12], we know that $\ell(A)=2\ln\sigma(A).$ For an element $A\in\mathrm{PU}(2,1)$, we take a representative $\tilde{A}\in\mathrm{U}(2,1)$ and define
[…]
Note that any two lifts of $A$ to $\mathrm{U}(2,1)$ differ by multiplication with a scalar $\lambda$ such that $|\lambda|=1$. So they have the same spectral radius. Consequently, Bergman translation length is well-defined on $\mathrm{PU}(2,1)$.
@@
for all $\gamma\in\pi_{1}(S_{g,k})$. The domination is said to be strict if $\ell(\rho_{1}(\gamma))<\ell(\rho_{2}(\gamma))$ for all $\gamma\in\pi_{1}(S_{g,k})$.
Let $\mathrm{PO}(2,1)$ be the identity component of the isometry group of ${\bf H}^{2}_{\mathbb{R}}$. Our main result is the following.
@@
Let $S_{g,k}$ be an oriented surface of negative Euler characteristic with at least one puncture. For a $T$-bent representation $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PU}(2,1)$, there exists a discrete and faithful representation $\rho_{0}:\pi_{1}(S_{g,k})\rightarrow\mathrm{PO}(2,1)$ that dominates $\rho$. Moreover, if $\gamma\in\pi_{1}(S_{g,k})$ is a peripheral loop, then the Bergman translation length of $\rho(\gamma)$ remains unchanged.
This is partial generalization of Deroin-Tholozan’s work in [DT15] for the punctured surfaces. We have used the $\tt Z$}-invariant (see Section 2.3) associated with a pair of adjacent real ideal triangles to establish our result. These invariants also motivate a notion of bending fiber for $T$-bent representations into $\mathrm{PU}(2,1)$, analogous to the construction in [BG25, Definition 2.14] for $\mathrm{PSL}_{n}(\mathbb{C})$ using Fock-Goncharov coordinates. In our context, we adopt the following version:
Definition 1.4 (Bending fiber for $\mathrm{PU}(2,1)$-representations).
Two $T$-bent representations $\rho_{1},\rho_{2}:\pi_{1}(S_{g,k})\to\mathrm{PU}(2,1)$ are said to lie in the same bending fiber if the absolute values of the $\tt Z$}-invariants assigned to each edge of $T$ coincide. In that case, we say that $\rho_{1}$ and $\rho_{2}$ are related by a bending deformation..
With this terminology, we can restate 1.3 in the language of bending fibers.
@@
The trace of an element $A\in\mathrm{SU}(2,1)$ also determines its geometric action and is directly associated with $\ell(A)$, as described in [Par12]. As discussed earlier, any two lifts of an element in $\mathrm{PU}(2,1)$ differ by a unitary scalar. While the trace of a lift $\tilde{A}$ may vary under such scalar multiplication, its modulus remains unchanged. Therefore, $|\mathrm{tr}(\rho(\gamma))|$ is a well-defined function on $\mathrm{PU}(2,1)$. As a corollary of 1.3, we also obtain:
@@
$|\mathrm{tr}(\rho(\gamma))|\leq|\mathrm{tr}(\rho_{0}(\gamma))|$
@@
We have also experimentally observed that the inequality does not extend for the discriminator function $f(\mathrm{tr}(\rho(\gamma)))$ (see 2.1). Further details are discussed in the Appendix.
Remark 1.
Note that 1.3 does not guarantee the existence of a representation $\rho_{0}:\pi_{1}(S_{g,k})\to\mathrm{PO}(2,1)$ that strictly dominates $\rho$. For instance, if a non-trivial, non-peripheral curve $\gamma\in\pi_{1}(S_{g,k})$ passes through a subsurface of $S_{g,k}$ where the $\tt Z$}-invariants are already real and positive, we get $\rho_{0}(\gamma)=\rho(\gamma)$.
@@ Acknowledgements
The authors would like to thank Subhojoy Gupta for indicating this problem to us and for motivations and useful discussions. The authors also thank Fanny Kassel and John R. Parker for valuable comments and discussions.
@@ 2.1. Complex hyperbolic plane and its isometries
Let us consider the non-degenerate, indefinite Hermitian form on $\mathbb{C}^{3}$ of signature (2,1) given by:
(2.1) $\langle\textbf{z},\textbf{w}\rangle=z_{1}\bar{w}_{3}+z_{2}\bar{w}_{2}+z_{3}\bar{w}_{1}$
where $\textbf{z}=(z_{1},z_{2},z_{3})^{t}$ and $\textbf{w}=(w_{1},w_{2},w_{3})^{t}\in\mathbb{C}^{3}$. In the canonical basis of $\mathbb{C}^{3}$, this Hermitian form is given by the matrix
$J=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$,
so that $\langle\textbf{z},\textbf{w}\rangle=\textbf{z}^{*}J\textbf{w}$.
Let $V_{-}$ and $V_{0}$ denote the set of negative and null vectors in $\mathbb{C}^{2,1}$ respectively, i.e.,
$V_{-}=\{\textbf{v}\in\mathbb{C}^{2,1}\ |\ \langle\textbf{v},\textbf{v}\rangle<0\}\text{ and }V_{0}=\{\textbf{v}\in\mathbb{C}^{2,1}\ |\ \langle\textbf{v},\textbf{v}\rangle=0\}.$
Note that both $V_{-}$ and $V_{0}$ are invariant under scalar multiplication by nonzero complex numbers, i.e., if $\textbf{v}\in V_{-}$ (respectively, $V_{0}$), then $c\textbf{v}\in V_{-}$ (respectively, $V_{0}$) for all $c\in\mathbb{C}^{*}$.
The complex hyperbolic 2-space $\mathbb{H}^{2}_{\mathbb{C}}$ is defined as the projectivization of $V_{-}$:
${\mathbf{H}}^{2}_{\mathbb{C}}:=\mathbb{P}V_{-}.$
Its boundary is given by the projectivization of null vectors: $\partial{\mathbf{H}}^{2}_{\mathbb{C}}=\mathbb{P}V_{0}$. This realization of ${\mathbf{H}}^{2}_{\mathbb{C}}$ corresponds to the Siegel model (see [Gol99, Chapter 4]), which is particularly convenient for describing the boundary via Heisenberg coordinates.
The Heisenberg coordinates (see [Par04, §4] or [Wil12, §2] for more details) of a point $(z_{1},z_{2},1)^{t}\in\partial{\mathbf{H}}^{2}_{\mathbb{C}}$ is given by $\zeta,t]$, where $z_{2}=\zeta\sqrt{2}$ and $z_{1}=-|\zeta|^{2}+it$.
@@
The holomorphic isometry group of ${\mathbf{H}}^{2}_{\mathbb{C}}$ is the projective unitary group $\mathrm{PU}(2,1)$, consisting of projective linear transformations that preserve the Hermitian form (up to scalar multiple). Explicitly, $\mathrm{PU}(2,1)=\mathrm{U}(2,1)/\{\lambda I\}$, where $\mathrm{U}(2,1)$ is the group of matrices $A\in\mathrm{GL}_{3}(\mathbb{C})$ satisfying $A^{*}JA=J$, with $J$ being the Hermitian matrix corresponding to the chosen form and $\lambda\in\mathbb{C}^{*}$. This group acts transitively and holomorphically on ${\mathbf{H}}^{2}_{\mathbb{C}}$, preserving both the complex structure and the Bergman metric. The isometry group Isom(${\mathbf{H}}^{2}_{\mathbb{C}}$) is generated by $\mathrm{PU}(2,1)$ and the antiholomorphic reflections along real Lagrangian planes (see section 2.2).
@@
[Gol99] Let $f(z)=|z|^{4}-8\Re(z^{3})+18|z|^{2}-27$. For $A\in\mathrm{SU}(2,1)$, the following holds:
  • AA has an eigenvalue $\lambda$ with $|\lambda|\neq 1$ if and only if $f(\mathrm{tr}(A))>0$; in this case, $A$ is loxodromic.
  • AA has a repeated eigenvalue if and only if $f(\mathrm{tr}(A))=0$; in this case, $A$ is either parabolic or special elliptic (i.e. elliptic with a repeated eigenvalue).
  • AA has distinct eigenvalues all of unit modulus if and only if $f(\mathrm{tr}(A))<0$; in this case, $A$ is elliptic.
@@ 2.2. Totally geodesic subspaces and ideal triangles
  • Complex lines are the intersections of ${\mathbf{H}}^{2}_{\mathbb{C}}$ with complex projective lines whenever these intersections are non-empty. They are biholomorphic to the Poincaré disk, and thus form embedded copies of ${\mathbf{H}}^{1}_{\mathbb{C}}$ inside ${\mathbf{H}}^{2}_{\mathbb{C}}$.
  • Totally real Lagrangian planes (also called real planes) are the intersections of ${\mathbf{H}}^{2}_{\mathbb{C}}$ with totally real 2-planes in $\mathbb{C}^{3}$, whenever these intersections are non-empty. They are isometric to the Klein–Beltrami model of the real hyperbolic plane $\mathbb{H}^{2}_{\mathbb{R}}$. Each such plane determines an antiholomorphic involution of ${\mathbf{H}}^{2}_{\mathbb{C}}$ fixing it pointwise; this involution is called its real symmetry.
@@
[Gol99, §7] Let $x=(x_{1},x_{2},x_{3})$ and $y=(y_{1},y_{2},y_{3})$ be two ideal triangles. Then there is a holomorphic (resp. antiholomorphic) isometry $A\in$ Isom $({\mathbf{H}}^{2}_{\mathbb{C}})$ with $A(x_{i})=y_{i}$ if and only if they have same (resp. opposite) Cartan invariant. Moreover, an ideal triangle has Cartan invariant 0 (resp. $\pm\pi/2$) if and only if it is contained in a real plane (resp. complex line).